# Zhang, Liu, and Zhao

### v4

@@ 4.3 Product placement problem
$\overline{Y}^{\prime}(X)=k\alpha_{1}\sin(\alpha_{2}X_{1})+k\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
where the sine function represents the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents the nonlinear trend with respect to $X_{2}$. We set $k=1$ unless otherwise specified. We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial, and we report performance of methods over 20 trials for each experimental setting. The sof method from Kallus and Mao (2023) is not applied for this two-stage SP, as it would require a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
@@
We first study the effect of the feature dimension $p$ on the performance of PADR under sparse and dense demand models similar to the discussion in Section 4.1.2. The results are reported in Figure 1p. Similar to the results in Section 4.1.2 for the Newsvendor problem, with sparse demand model, PADR performs well in low-dimensional cases but deteriorates as $p$ increases which may necessitate the proper parameter regularization. In contrast, forest-based prescriptive methods are robust to redundant features. With dense demand model, PADR and PO-PA have the best performance, with the former one slightly outperforming when the feature dimension increases to 50. Despite the increased complexity and the challenge in accurately approximating the true model, PADR continues to demonstrate superior performance for instances with dense demand model, high nonlinearity, and sparse demand models with small feature dimensions.
@@
(o) Sparse demand model
(p) Dense demand model
@@
We also study the effect of nonlinearity in the demand model by varying the coefficient $k$. Notice that the nonlinear structure of the demand model here is different from the one in Section 4.1 with the sine component function. The results for $p=20$ are provided in Figure 1r. The increasing gap in Figure 1q indicates that except for PADR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases. Figure 1r indicates that PADR is slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $k$ increases.
@@
Product placement problem: varying nonlinearity level $k$ (dense model, $p=20$) In (b), with the cost of the cart scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
@@ 5 Conclusion
We propose the PADR-based ERM method together with the ESMM algorithm to solve a broad class of SP problems with covariate information. For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems. To solve the ERM problem with a composite DC structure, we develop the ESMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling. We provide the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis. Numerical experiments indicate the advantages of our method over PO methods and prominent DR-based methods for various instances regarding the test cost, computational time, and robustness to nonlinearity of the underlying dependency. In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information. Extensions of the PADR-based ERM framework for more complicated SP problems are deserved to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
@@ 6 Proofs of Theoretical Consistency Results in Section 2
$f_{\widehat{\mathcal{C}}}(x)\coloneqq$ $\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}-\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},$ (20)
$\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(\widehat{y}-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$ $=\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}C\left\lVert\widehat{x}-\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}L_{0}\left\lVert\widehat{x}-\widehat{y}\right\rVert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\left\lvert\bar{f}_{0}(\widehat{x})-\bar{f}_{0}(\widehat{y})\right\rvert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y}),$
where the first inequality holds since $C=L_{0}/d_{\min}\geq L_{0}/\left\lVert\widehat{x}-\widehat{y}\right\rVert$. Moreover, this upper bound is attained by taking $\widehat{x}=\widehat{y}$ in the max operator. Similarly, the value of the second max-affine component of $f_{\widehat{\mathcal{C}}}$ at $\widehat{y}$ is $\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{y})$. Hence, we obtain $f_{\widehat{\mathcal{C}}}(\widehat{y})=\bar{f}_{0}(\widehat{y})$, verifying the interpolation. With $K\coloneqq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert$, for any $\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$, we obtain $\left\lVert C\widehat{x}\right\rVert_{\infty}\leq\frac{L_{0}\sqrt{p}}{2\epsilon}\bar{X}\leq\frac{L_{0}}{2}K^{1/p}$ and
@@
$\begin{array}[]{ll}&\mathcal{A}_{1}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},\\ &\mathcal{A}_{2}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}.\end{array}$
@@
$C\widehat{y}^{\top}(x-\widehat{y})+\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y})$ $=\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}-C\left\lVert x-\widehat{y}\right\rVert^{2}+\bar{f}_{0}(\widehat{y})\right]$
@@
$=C\widehat{x}^{\top}(x-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x}),$
where the second inequality is derived by the strict monotonicity of the quadratic function $t\mapsto L_{0}t-\frac{\sqrt{p}L_{0}}{2\epsilon}t^{2}$ on $[\epsilon/\sqrt{p},+\infty)$. The above result shows $\left\lVert\widehat{x}_{1}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$. Similarly, we have $\left\lVert\widehat{x}_{2}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{2}^{*}\in\mathcal{A}_{2}(x)$. It follows that
@@
^$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
@@
Lemma 6.3
Let $\widehat{\Phi}\coloneqq\{\phi(\cdot;\theta):\mathcal{U}\rightarrow\mathbb{R}\mid\theta\in\Theta\}$ be a class of functions parameterized by vectors in a compact set $\Theta\subseteq\mathbb{R}^{q}$ with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Suppose that there exist $M_{\widehat{\Phi}}>0$ and $L_{\widehat{\Phi}}>0$ such that $\max_{\theta\in\Theta}\left\lVert\phi(\cdot;\theta)\right\rVert_{\infty}\leq M_{\widehat{\Phi}}$ and $\left\lVert\phi(\cdot;\theta_{1})-\phi(\cdot;\theta_{2})\right\rVert_{\infty}\leq L_{\widehat{\Phi}}\left\lVert\theta_{1}-\theta_{2}\right\rVert$ for any $\theta_{1},\theta_{2}\in\Theta$. Then
@@
[…]
^$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
@@
[…]
Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$. Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $\Xi_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
@@
$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (21)
@@
$\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{4\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}.$ (22)
Now we prove the uniform generalization bound. By Lemma 6.5 and the inequality (21), with probability at least $21-\delta/$ 2,
$\max_{\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}}\left(\mathbb{E}_{X,Y}\phi(X,Y)-\frac{1}{n}\sum_{s=1}^{n}\phi(x^{s},y^{s})\right)\leq 2{\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
@@
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$

### v5

@@ 4.3 Product placement problem
$\overline{Y}^{\prime}(X)=\omega\alpha_{1}\sin(\alpha_{2}X_{1})+\omega\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
where the sine function represents the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents the nonlinear trend with respect to $X_{2}$. We set $\omega=1$ unless otherwise specified. We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial and report performance of methods over 10 trials for each experimental setting. For NN-DR, incorporating the second-stage problem into the loss function within the standard training framework is nontrivial. To address this difficulty, we treat the second-stage decision variables as additional model parameters and jointly optimize them along with the NN-DR parameters. It is worth noting that this implementation is tailored specifically to the problem at hand, and it underscores the limitation of NN–based approaches for multi-stage stochastic programs. The SOF method of Kallus and Mao (2023) is not included in this experiment, as applying it to a two-stage stochastic program would require substantial extensions involving second-order perturbation analysis, which are beyond the scope of this paper.
@@
We first study the effect of the feature dimension $p$ on the performance of PADR under sparse and dense demand models similar to the discussion in Section 4.1.2. The results are reported in Figure 1t. Similar to the results in Section 4.1.2 for the Newsvendor problem, with sparse demand model, PADR performs well in low-dimensional cases but deteriorates as $p$ increases which may necessitate the proper parameter regularization. NN-DR performs poorly in the sparse case, also suffering from overfitting. In contrast, forest-based prescriptive methods are robust to redundant features. With dense demand model, PADR and NN-DR have the best performance, with the former one slightly outperforming when $p$ is small. We also report the running time results in Figure 1v. PADR requires substantially more computational time in this problem, as the underlying optimization becomes more complex and involves multiple decision variables. In the current implementation, the decision rule is constructed separately for each dimension. Given so, we believe the computational efficiency of PADR could be further improved through vectorized multi-dimensional decision rule implementation and engineering optimizations in the ESMM surrogate construction.
Despite the increased complexity and the challenge in accurately approximating the true model, PADR continues to demonstrate superior performance for instances with dense demand model, high nonlinearity, and sparse demand models with small feature dimensions.
@@
(s) Sparse demand model
(t) Dense demand model
@@
[…]
(u) Sparse demand model
(v) Dense demand model
Running time in product placement problem with varying feature dimensions
@@
We also study the effect of nonlinearity in the demand model by varying the coefficient $\omega$. Notice that the nonlinear structure of the demand model here is different from the one in Section 4.1 with the sine component function. The results for $p=20$ are provided in Figure 1x. The increasing gap in Figure 1w indicates that except for PADR, NN-DR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases. Figure 1x indicates that PADR and NN-DR are slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $\omega$ increases.
@@
Product placement problem: varying nonlinearity level $\omega$ (dense model, $p=20$) In (b), with the cost of the CART scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
@@ 5 Conclusion
We propose the PADR-based ERM method together with the ESMM algorithm to solve a broad class of SP problems with covariate information. For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems. To solve the ERM problem with a composite DC structure, we develop the ESMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling. We provide the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis. Numerical experiments indicate the advantages of our method over PO methods and prominent DR-based methods for various instances regarding the test cost, computational time, and robustness to nonlinearity of the underlying dependency. In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information. Extensions of the PADR-based ERM framework for more complicated SP problems deserve to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
[…]
@@ 6 Proofs of Theoretical Consistency Results in Section 2
Proof of Proposition 2.1. It suffices to justify the bound when $d=1$. For $\epsilon>0$ satisfying $(\sqrt{p}\bar{X}/\epsilon)\geq 1$, we first construct a $(2\epsilon/\sqrt{p})$-grid of the cube $\bar{\mathcal{X}}\coloneqq[-\bar{X},\bar{X}]^{p}$, which is also an $\epsilon\text{-cover}$ of $\bar{\mathcal{X}}$ with $\ell_{2}$-norm, denoted by $\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$. Let $d_{\min}\coloneqq\underset{\widehat{x}_{1},\widehat{x}_{2}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}}),\,\widehat{x}_{1}\neq\widehat{x}_{2}}{\min}\left\lVert\widehat{x}_{1}-\widehat{x}_{2}\right\rVert=\frac{2\epsilon}{\sqrt{p}}$. The cardinality of $\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ satisfies
Given $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$, by McShane (1934, Theorem 1), we can extend $f_{0}$ to an $L_{0}$-Lipschitz function $\bar{f}_{0}:\bar{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying $\left\lVert\bar{f}_{0}\right\rVert_{\infty}\leq M_{0}$ and $\bar{f}_{0}(x)=f_{0}(x)$ for any $x\in\mathcal{X}$. Next, we construct a piecewise affine interpolation function to a set of points $\{(\widehat{x},\bar{f}_{0}(\widehat{x}))\mid\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
where the second inequality is derived by the strict monotonicity of the quadratic function $t\mapsto L_{0}t-\frac{\sqrt{p}L_{0}}{2\epsilon}t^{2}$ on $\epsilon/\sqrt{p},+\infty)$. The above result shows $\left\lVert\widehat{x}_{1}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$. Similarly, we have $\left\lVert\widehat{x}_{2}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{2}^{*}\in\mathcal{A}_{2}(x)$. It follows that
@@
$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
@@
It captures the richness of the class $\Phi$ by measuring the degree to which a hypothesis set can fit random noise. The following lemma quantifies the Rademacher complexity of a class of parametric functions, which is generalized from Ermoliev and Norkin (2013, Lemma B.2) with minor modifications made to accommodate the developments in this paper.
Lemma 6.3
@@
Proof of Lemma 6.3 Given a sample set $\Xi_{N}=\{u^{s}\}_{s=1}^{N}$ of size $N$, following the proof of Ermoliev and Norkin (2013, Lemma B.2), we derive that for any $\epsilon\in(0,2\mu]$,
$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
@@
Lemma 6.5 (Mohri et al. (2018), Theorem 3.3)
Let $\Phi$ be a function class mapping from $\mathcal{U}$ to [M,M][-M,M]. Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $\Xi_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
@@
$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{p,d}(\mu).$ (21)
@@
$\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{4\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}.$ (22)
Now we prove the uniform generalization bound. By Lemma 6.5 and the inequality (21), with probability at least $1-\delta/2$,
$\max_{\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}}\left(\mathbb{E}_{X,Y}\phi(X,Y)-\frac{1}{n}\sum_{s=1}^{n}\phi(x^{s},y^{s})\right)\leq 2{\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
@@
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq$ $\ \frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}$
$\ +2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$